Shift report 20030507 08:00-16:00

From: Apache (apache@pii3.brahms.bnl.gov)
Date: Wed May 07 2003 - 15:46:00 EDT

  • Next message: Apache: "Shift report 20030507 16:00-24:00"
    ------------------------------------------------------------
    Date: Wed May  7 15:46:00 2003, Shift: 08:00-16:00
    Supervisor: JJG, with: 
    Log book#14 pages 99 - 101
    
    Shift Summary:
    ------------------------------------------------------------
    1. Beam since 10.20 am : a miracle! 
    Inel rates from 5k to 2.5k. 
    Varying degrees of polarization (at end of shift no yellow polarization left).
    Beam dumped at 15.35 (for) 1 hr Phenix access.
    2. 
    -FS 4deg 1/4A: runs 8850 and 8851. With previous 8849: 100k trigger 6 in total.
    -FS 4deg 1/2A: runs 8852 and 8853  about 97k trigger 6.
    -FS 4 deg 1/1A: run 8854, about 15k trigger 6, 300 trigger 2.
    -FS 12 deg 1/4A: run 8855, 5.5 k trigger 6, 562 trigger 2.
    3. No detector problems.    
    ------------------------------------------------------------
    


    This archive was generated by hypermail 2.1.5 : Wed May 07 2003 - 15:50:41 EDT